Which equation is an identity?
a) 3w+6-w=4w-2(w-3)
b) 8-(5v+4)=-5v-4
c) 5y+7=5y-7
d) 4m-6=5m+9-m

Answer by ptaylor(2198)

The answer is a).
a)
3w+6-w=4w-2(w-3) get rid of parens
3w+6-w=4w-2w+6 collect like terms on each side of the equation
2w+6=2w+6---------------this is an identity---true for all values of w
b)
8-(5v+4)=-5v-4 get rid of parens
8-5v-4=-5v-4 collect like terms on each side of the equation
-5v+4=-5v-4
This problem does not have a solution. There are no values for v that will satisfy this equation.
c)
5y+7=5y-7
This problem does not have a solution. There are no values for y that will satisfy this equation.
d)
4m-6=5m+9-m collect like terms on each side of the equation
4m-6=4m+9
This problem does not have a solution. There are no values for m that will satisfy this equation.
Hope this helps---ptaylor